# ANote[fnsymbol]

### v2

@@ Abstract
Human beings perceive the real world through a spectrum of sensory modalities, encompassing auditory, visual, and linguistic faculties. The journey towards achieving Artificial General Intelligence (AGI) necessitates the development of models that can emulate these multifaceted perceptual capabilities and comprehensively understand these diversified data. To this end, we introduce Nexus-O, an industry-level omni-perceptive and -interactive model capable of efficiently processing Audio, Image, Video, and Text data in any combination and output audio/text in an end-to-end way. We systematically investigate Nexus-O by addressing three key research questions: First, how can models be efficiently designed and trained to achieve tri-modal alignment, understanding and reasoning capabilities across multiple modalities? Second, what approaches can be implemented to evaluate tri-modal model robustness, ensuring reliable performance and applicability in real-world scenarios? Third, what strategies can be employed to curate and obtain high-quality, real-life scenario speech datasets? For the first question, we design and pre-train Nexus-O based on the vision-language model, rather than the language model. By pre-training the model over high-quality synthetic audio data, our model is capable of tri-modal perception and interaction. For the second question, we introduce a new audio testbed, Nexus-O-audio, comprising diverse Automatic Speech Recognition (ASR) samples, spanning various real-world scenarios, such as corporate meetings and live stream. For the third question, we design the speech data synthesis pipeline to obtain high-quality speech training datasets, covering various real-world scenarios. Comprehensive experimentation and an in-depth analysis of tri-modal alignment over latent space demonstrate the advantages of our model on downstream tasks.
@@ 1 Introduction
Figure 1: An overview of the model’s performance. Nexus-O is capable of handling a wide range of tasks, demonstrating strong performance across various benchmarks. For Nexus-O-audio (EN) and (CN), the scores are expressed as reciprocal (1/scores).
@@
Since the mid-20th century, significant efforts have been devoted to designing systems of the capability to simulate human cognition, with the ultimate goal of developing Artificial General Intelligence (AGI) (Feng et al., 2024; Morris et al., 2023). Notable algorithms include Rule-based systems, Neuro-symbolic AI systems (Sheth et al., 2023), Bayesian nets systems (Heckerman, 2022), Neural networks systems (Hinton, 2006), etc. Among them, neural networks systems have emerged as a cornerstone thanks to their remarkable generalization capabilities. In this context, various neural architectures have been proposed, such as CNN (LeCun et al., 1995) and Transformer (Vaswani, 2017). The discovery of neural scaling laws (Kaplan et al., 2020) has further advanced the field, paving the way for Transformer-based Large Language Models (LLMs), such as ChatGPT (Kaplan et al., 2020), which centre on language as the medium for perceiving and interacting with both human society and the physical world. By aligning multiple modalities around the language, Multimodal Large Language Models (MLLMs) have gained widespread recognition as a promising pathway toward AGI (Liang et al., 2024; Zhang et al., 2024a; NVIDIA et al., 2025; Ma et al., 2025; Hong et al., 2025).
@@
MLLMs, involving language, vision, and auditory modalities, can generally be divided into three categories: vision-language model (Wang et al., 2024), audio-language model (Chu et al., 2024), and vision-language-audio model (Fu et al., 2025). While alignment among three modalities offers a more comprehensive simulation of human cognitive processes than two-modality alignment, it presents significantly greater challenges. These challenges include limited accessibility to tri-modal datasets, higher computational resource requirements, and the complexity of aligning features across three modalities. These limitations hinder the advancement of omni-perceptive and interactive AGI systems. To address these challenges, we explores the development of a fully end-to-end tri-modal system capable of omni-perception and seamless interaction with the world in this work.
To achieve this goal, several key research questions need to be addressed. First, what methodologies and strategies can be employed to curate and obtain high-quality, real-scenario speech datasets? Second, how can models be efficiently designed and trained to achieve tri-modal alignment, understanding and reasoning capabilities across multiple modalities? Third, what approaches can be implemented to evaluate audio-modal robustness, ensuring reliable performance and applicability in real-world scenarios?
In response to these questions, we propose Nexus-O with the following three approaches: (1) We systematically design the speech data synthesis pipeline to obtain high-quality speech training datasets, spanning various real-world scenarios, such as corporate meetings, live broadcasting scenarios, etc. (2) We start by designing and pre-training Nexus-O based on the vision-language model, especially Qwen2.5-VL (Bai et al., 2025) [^1] In this report, Qwen2.5-VL refers to the instruction version., rather than a language model, as is typically chosen in most omni models (Fu et al., 2024c; Xie & Wu, 2024a; b; Li et al., 2024a). By pre-training the model over high-quality synthetic audio data, Nexus-O is capable of tri-modal perception and interaction. This strategy offers high training efficiency while reducing the computational resource burden derived from the training vision-language alignment stage, which is an efficient and effective option for most institutions with low computational resources. (3) For a fair evaluation of the model in real-world scenarios, we introduce a new audio testbed, i.e., Nexus-O-audio, comprising diverse Automatic Speech Recognition (ASR) samples, spanning various real-world scenarios.
@@
The experimental results demonstrate the following five key findings: (1) In the visual understanding task, Nexus-O exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, indicating the effectiveness of our training strategy. (2) Within the Audio English QA benchmark, the model achieves better accuracy than the same-period competitor (i.e, MiniCPM-o2.6-7B) in the LLaMA Q. benchmark (Nachmani et al., 2024). (3) In our real-world ASR benchmarks, Nexus-O achieves outstanding performance, which underscores its robustness in real scenarios. (4) In the Speech-to-Text Translation benchmark, our model outperforms Qwen2-Audio-Instruct-7B. (5) Furthermore, an in-depth analysis of tri-modal alignment reveals that incorporating the audio modality enhances representational alignment. Even in the absence of audio input, our model delivers performance comparable to the backbone model. The main contributions of this paper are summarized as follows:
@@
  • We systematically design the speech data synthesis pipeline to obtain high-quality speech datasets to facilitate the rapid deployment of the model in real-world scenarios.
  • We propose an omni-perceive and -interactive tri-modal model, i.e., Nexus-O, which supports any combination of audio, image/video, and text inputs. Different from existing approaches that treat audio as an auxiliary, our model enables a joint understanding of these modalities.
  • To evaluate the performance of the tri-modal model, we propose the Nexus-O-audio testbed, covering various real-world scenarios, such as corporate meetings, live broadcasting scenarios, etc. We intend to release the Nexus-O-audio testbed to establish a new standard for fostering innovation in addressing complex real-world applications.
@@
  • We comprehensively evaluate the performance of Nexus-O on various benchmarks and conduct an in-depth analysis of tri-modal alignment from the perspective of representation space. Experimental results illustrate the efficiency and the effectiveness of our tri-modal model.
The structure of this work is organised as follows: In Section 2, we first identify the research gap between our model and other contemporaneous approaches, and then discuss the limitations of current speech benchmarks. In Section 3, we present our tri-modal model architecture, the composition of training data and audio data synthesis pipeline, pre-training strategies, and the Nexus-O-audio testbed. In Section 4, we first address the alignment of downstream tasks, including vision-language evaluation and audio-language evaluation, and then evaluate representational alignment. Finally, we conclude with a summary of our findings and outline our future research directions in Section 5.
@@ 2.1 MLLMs
Large Language Models (LLMs) have demonstrated remarkable capabilities in understanding and reasoning over textual knowledge (Touvron et al., 2023a; b; Dubey et al., 2024; Yang et al., 2024a; Jiang et al., 2023). Building on these advancements, a bunch of studies (Chen et al., 2024a; b; Dai et al., 2022) have extended the understanding and alignment capabilities of LLMs to visual knowledge, addressing challenges such as vision-language alignment and instruction-following, by introducing Multimodal Large Language Models (MLLMs). Notable models include Qwen-VL (Wang et al., 2024; Yang et al., 2024a), InternVL (Chen et al., 2024c), etc. However, in complex visual reasoning tasks, such as video analysis, MLLMs typically struggle to perform well due to the heterogeneity among diverse modalities, where different modalities have distinct feature granularities, representations, and geometries in a shared latent space (Zhang et al., 2024e; c). To further improve their performance in visual understanding tasks, recent works have expanded MLLMs by incorporating audio capabilities, such as LLaMA-Omni (Fang et al., 2024a), VITA-1.0 (Fu et al., 2024c), Mini-Omni2 (Xie & Wu, 2024b), Baichuan-Omni (Li et al., 2024a), and MinMo (Chen et al., 2025a). However, these models exhibit certain limitations: LLaMA-Omni and MinMo lacks visual capabilities, Baichuan-Omni does not have end-to-end Text To Speech (TTS) capabilities, and VITA-1.0 is constrained by the limited capability of its backbone model.
Recent models, such as VITA-1.5 (Fu et al., 2025), MiniCPM-o2.6  [^2] https://github.com/OpenBMB/MiniCPM-o, Baichuan-omni-1.5333https://github.com/baichuan-inc/Baichuan-Omni-1.5, and OpenOmni (Luo et al., 2025), have been proposed and have achieved state-of-the-art performance. However, these are pretrained based on the LLMs rather than MLLMs. While pre-training from scratch can enhance multimodal alignment, this strategy significantly increases training time and resource demands due to the necessity of additional visual-language alignment, which relies on large-scale pre-training datasets. To improve training efficiency and to reduce resource burden, we start by pre-training over MLLM, specifically Qwen2.5-VL-7B, and introduce Nexus-O, an advanced unified tri-modal LLM designed to enable audio understanding and generation across arbitrary combinations of modalities. This strategy is friendly with institutions with limited computation resources.
@@ 2.2 Audio Modality Evaluation
With the rapid advancement of MLLMs, a wide range of multimodal benchmarks has been proposed and widely adopted for evaluation, ranging over vision (Xie et al., 2023; Gan et al., 2024; Fu et al., 2024b; Yue et al., 2024; Yang et al., 2025), audio (Conneau et al., 2023; Yang et al., 2024b; Fu et al., 2024a), omni(Li et al., 2024b), etc. In the speech domain, existing benchmarks, e.g., Fleurs (Conneau et al., 2023), Aishell2 (Du et al., 2018a), LibriSpeech (Pratap et al., 2020b), and Common Voice (Ardila et al., 2019), provide linguistically diverse datasets under varying acoustic conditions, facilitating Automatic Speech Recognition (ASR) evaluation. Additionally, AIR-Bench (Yang et al., 2024b) introduces a pioneering generative evaluation framework encompassing speech, natural sounds, and music, employing novel audio mixing strategies and GPT-4 for unified, objective, and reproducible assessments.
However, the assessment of speech models should be closely aligned with real-world environments to ensure that the evaluation results faithfully represent the demands and requirements of practical and industrial applications. We observe that these existing speech benchmarks present three limitations. First, the small data scale of these benchmarks typically results in significant performance variance. Second, the benchmarks primarily focus on controlled scenarios, making it difficult to evaluate models in dynamic and complex real-world settings, e.g., meetings or live broadcast environments. Third, scores from the existing ASR benchmarks generally fail to align closely with the actual user interaction experience, thereby limiting their practical utility. To address these limitations, we propose Nexus-O-audio testbed, comprising 8.1k and 2.8k of Chinese and English ASR samples, spanning various application domains, such as corporate meetings and live broadcasting scenarios.
In addition to assessing audio downstream task alignment, we conduct an in-depth evaluation of whether audio modality can enhance vision-language alignment, as well. Intuitively, the audio modality can provide complementary information to the visual and language modalities, enhancing mutual information and feature alignment, thereby improving the performance on visual understanding tasks (Song et al., 2023; Choi et al., 2024; Zhang et al., 2024b).
Figure 2: Architecture of Nexus-O, which is designed to accept any combination of input modalities and generates output in either the language or audio modality, where the Auto-Regressive (AR) audio decoder takes a special start token embedding and the last language embedding as input to generate the hierarchical discrete audio codes in an auto-regressive manner. These codes are subsequently fed into a pretrained audio generator to produce the final audio output.
3 Nexus-O
In this section, we introduce the architecture of Nexus-O in Section 3.1, training data composition and audio data synthesis pipeline in Section 3.2, pre-training strategies in Section 3.3, and Nexus-O-audio testbed in Section 3.4, respectively.
3.1 Architecture
@@
Nexus-O utilises the language model component of Qwen2.5-VL-7B (Bai et al., 2025) as the foundational language model, consisting of 28 causal Transformer layers.
@@ 3.2 Training Data
Figure 3: Audio dataset synthesis pipeline. In the current version, our testbed only supports the ASR task. We will further incorporate various audio-modal tasks, including the AQA and AVQA tasks. Both are currently under development.
@@
As detailed in Table 1, the training data for Nexus-O encompasses a wide range of domains within speech processing. These domains include recognition tasks such as ASR and Speech Emotion Recognition (SER), reasoning tasks include Automatic Audio Caption and Speech Emotion Recognition, translation tasks like Speech-to-Text Translation (S2TT), generation tasks such as Speech Synthesis, and interactive tasks involving Speech-to-Speech Chat and Speech-to-Text Chat. The pretrained datasets include (1) public corpora (50%): Aishell1 (Bu et al., 2017), Wenetspeech (Zhang et al., 2022), Librispeech (Panayotov et al., 2015a), Kensho (O’Neill et al., 2021), PeopleSpeech (Galvez et al., 2021), MLS (Pratap et al., 2020a), and Gigast (Ye et al., 2023) and (2) real-life in-house data (50%).

### v3

@@ Abstract
Human beings perceive the real world through a spectrum of sensory modalities, encompassing auditory, visual, and linguistic faculties. This work proposes an industry-level omni-modal large language model (LLM) pipeline that integrates auditory, visual, and linguistic modalities to overcome challenges such as limited tri-modal datasets, high computational costs, and complex feature alignments. Our pipeline consists of three main components: First, a modular, end-to-end framework enabling flexible configuration of various encoder–LLM–decoder architectures. Second, a lightweight training strategy that pre-trains audio-language alignment on the state-of-the-art vision-language model Qwen2.5-VL, thus avoiding the costly pre-training of vision-specific modalities. Third, an audio synthesis pipeline that generates high-quality audio-text data from diverse real-world scenarios, supporting applications such as Automatic Speech Recognition and Speech-to-Speech chat. To this end, we introduce an industry-level omni-modal LLM, Nexus. Extensive experiments validate the efficacy of our pipeline, yielding the following key findings: (1) In the visual understanding task, Nexus exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, validating the efficiency of our training strategy. (2) Within the English Spoken Question-Answering task, the model achieves better accuracy than the same-period competitor (i.e, MiniCPM-o2.6-7B) in the LLaMA Q. benchmark. (3) In our real-world ASR testset, Nexus achieves outstanding performance, indicating its robustness in real scenarios. (4) In the Speech-to-Text Translation task, our model outperforms Qwen2-Audio-Instruct-7B. (5) In the Text-to-Speech task, based on pretrained vocoder (e.g., Fishspeech1.4 or CosyVoice2.0), Nexus is comparable to its backbone vocoder on Seed-TTS benchmark. (6) An in-depth analysis of tri-modal alignment reveals that incorporating the audio modality enhances representational alignment between vision and language.
@@ 1 Introduction
Figure 1: An overview of the model’s performance. Nexus is capable of handling a wide range of tasks, demonstrating strong performance across various benchmarks. For Nexus-audio (EN) and (CH), the scores are expressed as reciprocal (1/scores).
@@
Since the mid-20th century, significant efforts have been devoted to designing systems of the capability to simulate human cognition, with the ultimate goal of developing Artificial General Intelligence (AGI). Notable algorithms include Rule-based systems, Neuro-symbolic AI systems, Neural networks systems (Hinton, 2006), etc. Among them, neural networks systems have emerged as a cornerstone thanks to their remarkable generalization capabilities. In this context, various neural architectures have been proposed, such as CNN (LeCun et al., 1995) and Transformer (Vaswani, 2017). The discovery of neural scaling laws (Kaplan et al., 2020) has further advanced the field, paving the way for Transformer-based Large Language Models (LLMs), such as ChatGPT, which centre on language as the medium for perceiving and interacting with both human society and the physical world. By aligning multiple modalities around the language, Multimodal Large Language Models (MLLMs) have gained widespread recognition as a promising pathway toward AGI.
@@
MLLMs, involving language, vision, and auditory modalities, can generally be divided into three categories: vision-language model (Wang et al., 2024), audio-language model (Chu et al., 2024b), and vision-language-audio model (Fu et al., 2025). While alignment among three modalities offers a more comprehensive simulation of human cognitive processes than two-modality alignment, it presents significantly greater challenges. These challenges include limited accessibility to tri-modal datasets, higher computational resource requirements, and the complexity of aligning features across three modalities. These limitations hinder the advancement of omni-perceptive and interactive AGI systems. To address these challenges, we propose a comprehensive, industry-level omni-modal LLM pipeline that comprises a modularised end-to-end framework, a lightweight training strategy, and an audio data synthesis pipeline:
First, in the modularised framework, each modality (audio, image/video, and text) is processed using its own dedicated pretrained encoder. This design ensures that the inherent characteristics and fine-grained details unique to each modality are effectively preserved within their respective latent representations. Subsequently, these modality-specific representations are integrated and aligned within a LLM, and the resulting unified representations are subsequently supplied to decoupled decoders designed for each modality. This modular formalism design facilitates efficient and flexible configuration of various encoder–LLM–decoder architectures, thereby potentially contributing to the advancement of AGI by integrating components such as language models, world models, and embodied intelligence.
Second, most omni-modal models, such as Qwen2.5-omni (Xu et al., 2025), VITA1.5 (Fu et al., 2024c), etc., necessitate a preliminary pre-training stage for vision alignment, understanding, and reasoning. This stage typically demands substantial computational resources due to the processing of long sequences of image tokens, which is consequently less accessible to smaller research institutions and companies. To mitigate this limitation, we only pre-train audio-language alignment over the current state-of-the-art vision-language model, Qwen2.5-VL (Bai et al., 2025) [^1] In this report, Qwen2.5-VL refers to the instruction version.. Due to the inherent alignment between audio and text modalities, this training strategy effectively integrates audio understanding and generation capabilities while preserving the model’s established vision comprehension and reasoning functions.
Third, high-quality audio training data are often less accessible in both academy and industry. To address this limitation, we propose an audio synthesis pipeline that generates high-quality audio-text data spanning a broad spectrum of real-world scenarios, such as live streaming and dialogue, thereby supporting various downstream tasks such as Automatic Speech Recognition (ASR) and Speech-to-Speech chat. By incorporating real-world audio data at the pretraining stage, the model can be fast deployed for real-world applications. Based on our audio synthesis pipeline, we introduce a real-life ASR testset, named as Nexus-audio.
@@
To this end, we introduce Nexus222“Nexus”: connect the strongest models to form an even more powerful integrated model., an industry-level omni-modal LLM. Extensive experiments validate the efficacy of our pipeline, yielding the following key findings: (1) In the visual understanding task, Nexus exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, validating the efficiency of our training strategy. (2) Within the English Spoken Question-Answering task, the model achieves better accuracy than the same-period competitor (i.e, MiniCPM-o2.6-7B) in the LLaMA Q. benchmark (Nachmani et al., 2024). (3) In our real-world ASR testset, Nexus achieves outstanding performance, indicating its robustness in real scenarios. (4) In the Speech-to-Text Translation task, our model outperforms Qwen2-Audio-Instruct-7B. (5) In the Text-to-Speech task, based on pretrained vocoder (e.g., Fishspeech1.4 or CosyVoice2.0) Nexus is comparable to its backbone vocoder on Seed-TTS benchmark Anastassiou et al. (2024).
To further investigate the interplay among three modalities, we conducted an in-depth analysis of tri-modal alignment. Our findings indicate that (6) the inherent correspondence between audio and language can serve to facilitate and reinforce vision-language alignment, indicating the efficiency of our training strategy. The main contributions of this paper are summarised as follows:
  • We propose a comprehensive, industry-level omni-modal LLM pipeline that comprises a modularised end-to-end framework, a lightweight training strategy, and an audio data synthesis pipeline.
  • Based on the proposed pipeline, we introduce Nexus, an industry-level omni-modal LLM. It is designed to support any combination of audio, image/video, and text inputs, and it is capable of generating outputs in either audio or language modalities.
@@
  • We systematically design the speech data synthesis pipeline to obtain high-quality, real-life speech datasets, covering various real-world scenarios, such as corporate meetings, live broadcasting scenarios, etc., to facilitate the rapid deployment of the model in real-world scenarios. Moreover, we introduce the Nexus-audio testset to evaluate its robustness in real life ASR task.
@@
  • We comprehensively evaluate the performance of Nexus on different benchmarks and conduct an in-depth analysis of vision-language alignment by incorporating audio modality within the representation space. Experimental results indicate the merits of our proposed pipeline.
To the best of our knowledge, we are the first to propose the omni-modal LLM from the modularised design perspective that “connect the strongest models to form an even more powerful integrated model”. We intend to release the pretrained checkpoint, experimental pipeline, and test set, thereby establishing a new standard to foster innovation in addressing complex, real-world applications.
@@ 2.1 MLLMs
Large Language Models (LLMs) have demonstrated remarkable capabilities in understanding and reasoning over textual knowledge (Yang et al., 2024a). Building on these advancements, a bunch of studies (Dai et al., 2022) have extended the understanding and alignment capabilities of LLMs to visual knowledge, addressing challenges such as vision-language alignment and instruction-following, by introducing Multimodal Large Language Models (MLLMs). Notable models include Qwen-VL (Wang et al., 2024), InternVL (Chen et al., 2024), etc. However, in complex visual reasoning tasks, such as video analysis, MLLMs typically struggle to perform well due to the heterogeneity among diverse modalities, where different modalities have distinct feature granularities, representations, and geometries in a shared latent space (Zhang et al., 2024). To further improve their performance in visual understanding tasks, recent works have expanded MLLMs by incorporating audio capabilities, such as LLaMA-Omni (Fang et al., 2024a), VITA-1.0 (Fu et al., 2024c), Mini-Omni2 (Xie & Wu, 2024b), Baichuan-Omni (Li et al., 2024a), and MinMo (Chen et al., 2025a). However, these models exhibit certain limitations: LLaMA-Omni and MinMo lacks visual capabilities, Baichuan-Omni does not have end-to-end Text To Speech (TTS) capabilities, and VITA-1.0 is constrained by the limited capability of its backbone model.
Recent models, such as VITA-1.5 (Fu et al., 2025), MiniCPM-o2.6  [^3] https://github.com/OpenBMB/MiniCPM-o, Baichuan-omni-1.5444https://github.com/baichuan-inc/Baichuan-Omni-1.5, and OpenOmni (Luo et al., 2025), have been proposed and have achieved state-of-the-art performance. However, these models usually require a visual pretraining stage. While pretraining from scratch can enhance multimodal alignment, this strategy significantly increases training time and resource demands due to long sequence of image or video tokens and large pretraining vision data, which is less friendly with institutions with limited computation resources. To improve training efficiency, we start by pre-training over MLLM, specifically Qwen2.5-VL-7B. Owing to the natural alignment between audio and language modalities, the integration of the audio modality has minimal impact on the pretrained geometric and manifold structures underlying the vision-language alignment. As a result, the model’s established capabilities in vision comprehension and reasoning are effectively preserved.
@@ 2.2 Audio Modality Evaluation
With the rapid advancement of MLLMs, a wide range of multimodal benchmarks has been proposed and widely adopted for evaluation, ranging over vision (Fu et al., 2024b), audio (Fu et al., 2024a), omni (Li et al., 2024b), etc. In the speech domain, existing benchmarks, e.g., Fleurs (Conneau et al., 2023), Aishell2 (Du et al., 2018a), LibriSpeech (Pratap et al., 2020b), and Common Voice (Ardila et al., 2019), provide linguistically diverse datasets under varying acoustic conditions, facilitating Automatic Speech Recognition (ASR) evaluation. Additionally, AIR-Bench (Yang et al., 2024b) introduces a pioneering generative evaluation framework encompassing speech, natural sounds, and music, employing novel audio mixing strategies and GPT-4 for unified, objective, and reproducible assessments.
However, the assessment of speech models should be closely aligned with real-world environments to ensure that the evaluation results faithfully represent the demands and requirements of practical and industrial applications. We observe that these existing speech benchmarks present three limitations. First, the small data scale of these benchmarks typically results in significant performance variance. Second, the benchmarks primarily focus on controlled scenarios, making it difficult to evaluate models in dynamic and complex real-world settings, e.g., meetings or live broadcast environments. Third, quantitative evaluation scores from the existing ASR benchmarks generally fail to align closely with the actual user interaction experience, thereby limiting their practical utility. To address these limitations, we propose Nexus-audio testset, comprising 8.1k and 2.8k of Chinese and English ASR samples, spanning various application domains, such as corporate meetings and live broadcasting scenarios.
Figure 2: Modularised Architecture, which is designed to accept any combination of input modalities and generates output in either the language or audio modality, where the Auto-Regressive (AR) audio decoder takes a special start token embedding and the last language embedding as input to generate the hierarchical discrete audio codes in an auto-regressive manner. These codes are subsequently fed into a pretrained audio generator to produce the final audio output.
3 Nexus
In this section, we introduce the architecture of Nexus in Section 3.1, training data composition and audio data synthesis pipeline in Section 3.2, pre-training strategies in Section 3.3, respectively.
3.1 Modularised Architecture
@@
Nexus utilises the language model component of Qwen2.5-VL-7B (Bai et al., 2025) as the foundational language model.
@@ 3.2 Training Data
Figure 3: Audio dataset synthesis pipeline, the upper component represents the text-to-audio branch, while the lower component corresponds to the audio-to-text branch. Both components incorporate an equal proportion of in house, real-world samples.
@@
As detailed in Table 1, the training data for Nexus encompasses a wide range of domains within speech processing. These domains include (i) recognition tasks such as ASR and Speech Emotion Recognition (SER), (ii) reasoning tasks include Automatic Audio Caption and Speech Emotion Recognition, (iii) translation tasks like Speech-to-Text Translation (S2TT), generation tasks such as Speech Synthesis, and (iv) interactive tasks involving Speech-to-Speech Chat and Speech-to-Text Chat. The pretrained datasets include (1) public corpora (50%): Aishell1 (Bu et al., 2017), Wenetspeech (Zhang et al., 2022), Librispeech (Panayotov et al., 2015a), Kensho (O’Neill et al., 2021), PeopleSpeech (Galvez et al., 2021), MLS (Pratap et al., 2020a), and Gigast (Ye et al., 2023) and (2) real-life in-house data (50%).
Table 1 presents the statistical information for the audio training dataset. Unlike most Omni-MLLMs, which typically report duration in hours Chen et al. (2025a), our summary is based on the number of audio samples. The total duration of the dataset is approximately 30k hours, this is considerably smaller than the durations employed in other works, such as MinMo (pretrained over 1.4 million hours) Chen et al. (2025a) and Qwen2-Audio (pretrained over 520k hours) Chu et al. (2024a).